22 Suppose sin v = 3. Evaluate (a) cos u (c) tan u (b) sin u (d) cos v (e) tan v.
Answer #1

since = 3/7 cose= sin sinne = $ 1 - (3) 2 = √ 1 - 4 49 40 vo. 49 7 and tance - since colo 3/7 2/7500 3 arto